# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

### v1

@@ 2 The basic ingredients
Let us first see how the non-minimal term contributes to the gravitational interaction, without any $\Lambda$. The simplest process is given by the first of 1. From 6, we see that the non-minimal one graviton - two scalar vertex function reads : $i\xi\kappa(-q^{2})\eta_{\mu\nu}/2$. The $k_{2}-k^{\prime}_{2}$ vertex is the usual minimal one graviton - two scalar vertex. The Feynman amplitude reads,
@@
$i{\cal M}=-i.i\kappa\frac{i\xi\kappa(-q^{2})\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (10)
[…]
@@
Figure 1: The tree and the box diagrams for 2-2 scattering with graviton exchanges. Solid and wavy lines respectively denote scalar and graviton propagators. The thick circle denotes non-minimal vertex of ${\cal O}(\xi\kappa)$. The external momenta are on shell, ${k^{2}_{1}}={k^{\prime}}_{1}^{2}=-m_{1}^{2}$, ${k^{2}_{2}}={k^{\prime}}_{2}^{2}=-m_{2}^{2}$. None of the diagrams yields any non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to gravitational potential. Similar conclusion holds for the crossed box and other one loop diagrams in the absence of any $\Lambda$-three graviton vertex. See main text for detail.
@@
Figure 3: The vacuum polarisation diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. One of the cubic graviton vertex corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8. Also as earlier, the thick circle denotes non-minimal vertex.
@@ 3.1 The seagull
$-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-\frac{\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
[…]
$\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$=-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,$ (11)
@@
where $\mu$ is the mass scale for renormalisation. Likewise if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain vanishing contribution.
@@
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})},$ (12)
@@
If we place the non-minimal vertex on the $k_{2}$ vertex, we obtain exactly the same result as above. Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result, Thus we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,

### v2

@@ 2 The basic ingredients
Let us first see how the non-minimal term contributes to the gravitational interaction, without any $\Lambda$. The tree level process is given by the first of 1. From 6, we see that the non-minimal one graviton-two scalar vertex function reads: $i\xi\kappa(-q^{2})\eta_{\mu\nu}/2$. The $k_{2}-k^{\prime}_{2}$ vertex is taken to be the usual minimal one graviton-two scalar vertex. The Feynman amplitude reads,
@@
$i{\cal M}=-i.i\kappa\frac{i(-q^{2})\xi\kappa\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (11)
where $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$. The long range gravitational potential is defined in the non-relativistic limit by the Fourier transform of ${\cal M}/4m_{1}m_{2}$ in the first Born approximation, originating from the functions non-analytic in $q^{2}$. However, the potential corresponding 11 is proportional to $\delta^{3}(\vec{r})$, i.e., not a long range one we are interested in. It can be checked explicitly that analogous things occurs for the box diagram of 1. As a further explicit example, let us consider the triangle diagram generated by pinching the $k_{1}-l$-line of the box diagram. This will generate a two scalar-two graviton vertex in the upper half of the diagram. Let us first imagine that this vertex is non-minimal, produced by 6. It is not difficult to see that this diagram generates a momentum integral of the kind
$\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{l\cdot(l-q)\ k_{2}\cdot(l-k^{\prime}_{2})\ k^{\prime}_{2}\cdot(l-k^{\prime}_{2})}{l^{2}(l-q)^{2}\left((l-k^{\prime}_{2})^{2}+m_{2}^{2}\right)},$ (12)
which gives a leading potential $\sim G^{2}\xi m_{2}^{2}/m_{1}r^{4}$ in the non-relativistic limit. Note the appearance of the mass of the scalar in the denominator. Even though long range, this potential is much subleading and we wish to look for leading results than this. Similar result is found if instead we pinch the $(l+k_{2})$-line of the second of 1. If on the other hand, we take either the $k_{2}$ or the $k^{\prime}_{2}$ vertex to be non-minimal, we encounter integrals like
$\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{l^{2}\ k_{2}\cdot(l+k_{2})}{l^{2}(l-q)^{2}\left((l+k_{2})^{2}+m_{2}^{2}\right)}\quad{\rm Or,}\quad\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{(l-q)^{2}\ k^{\prime}_{2}\cdot(l+k_{2})}{l^{2}(l-q)^{2}\left((l+k_{2})^{2}+m_{2}^{2}\right)},$
both of which are independent of $q$ and hence makes no contribution to the long range potential. Likewise, for the other one loop diagrams as well, we find either they do not yield any potential at all, or their leading contribution is ${\cal O}(r^{-4})$.
As we have already emphasised, evidently the above situation arises due to the transfer momentum carried by the graviton appearing in the numerator. This motivates us to bring in the $\Lambda$-three graviton vertex mentioned above, and to see whether we may obtain contribution leading compared to $r^{-4}$. There can be three relevant diagrams involving this vertex at one loop, which we wish to evaluate below. 9 will be one of the chief ingredients for the following calculations.
@@
Figure 1: The tree and the box diagrams for 2-2 scattering with graviton exchanges. Solid and wavy lines respectively denote scalar and graviton propagators. The thick circle denotes a non-minimal vertex of ${\cal O}(\xi\kappa)$. The external momenta are on shell, ${k^{2}_{1}}={k^{\prime}}_{1}^{2}=-m_{1}^{2}$, ${k^{2}_{2}}={k^{\prime}}_{2}^{2}=-m_{2}^{2}$. Both diagrams yield no non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to the long range gravitational potential. The leading contribution behaves as $r^{-4}$ at ${\cal{O}}(\xi G^{2})$ from some other diagrams. See the main text for details.
@@
Figure 3: The vacuum polarisation diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. One of the cubic graviton vertices corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8. Also, as earlier, the thick circle denotes a non-minimal vertex.
@@ 3.1 The seagull
$\begin{split}&-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}\eta^{\mu^{\prime}\lambda^{\prime}}\eta^{\nu^{\prime}\rho^{\prime}}\eta_{\alpha\beta}P_{\mu\nu\mu^{\prime}\nu^{\prime}}P^{\alpha\beta\gamma\delta}P_{\lambda^{\prime}\rho^{\prime}}{}^{\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+\mu^{4\epsilon}\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ =&-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,\end{split}$ (13)
@@
where $\mu$ is the mass scale for renormalisation. Likewise, if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain a vanishing contribution.
@@
$\begin{split}=&-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\\ &+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\end{split}$
$\begin{split}=&-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\end{split}$ (14)
@@
If we place the non-minimal vertex on the $k_{2}$ vertex, we obtain exactly the same result as above. Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result. Thus, we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,
@@
By 10, the above amplitude will give an $r^{-1}$ potential in the non-relativistic limit. Note that there is no subleading contribution to the potential from the seagull diagrams. This is simply because only the second and the third diagrams of 3 (along with the pairwise interchange of the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices) contribute to the above potential. Due to the placement of the nonminimal vertex in these diagrams we have $(l-q)^{2}$ or $l^{2}$ appearing in the numerator of the corresponding scattering amplitudes. This results in the factorisation of $q^{-2}$ and momentum integral independent of the transfer momentum $q$, 14. Accordingly, the contribution in 16 is exact, and there is no other terms (like $q^{-1}$ or $\ln q^{2}$) which can make subleading contributions to the potential.